# A unified model for long-period radio transients and white dwarf binary pulsars

### v1

@@ Abstract
Long-period radio transients (LPTs) represent a recently uncovered class of Galactic radio sources exhibiting minute-to-hour periodicities and highly polarised pulses of second-to-minute duration. Their phenomenology does not fit exactly in any other class, although it might resemble that of radio magnetars or of white dwarf (WD) radio emitting binary systems. Notably, two LPTs with confirmed multi-wavelength counterparts have been identified as synchronised white dwarf–M dwarf binaries (polars). Meanwhile, systems such as AR Scorpii and J1912--44 exhibit short-period pulsations in a hr-tight orbit, with polarized radio emission proposed to be generated by the interaction of the WD magnetosphere with the low-mass companion wind.
Here we demonstrate that both LPTs and WD binary pulsars can be explained within a single geometric model in which radio emission is triggered when the magnetic pole of a rotating white dwarf intersects its companion’s wind in the binary orbital plane. We use a 36-year timing baseline to infer the orbital period and binary geometry solely from radio data of GPM J1839--10, the longest-active LPT known. The model naturally predicts its intermittent emission and double-pulse structure. Crucially, we show that the beat period between the spin and the orbit matches the observed pulse substructure and polarisation signatures, providing strong support for the model. Applying this model to the WD binary pulsar J1912--44  it can also reproduce the system emission and geometry. Our results place GPM J1839--10, and other LPTs in general and radio emitting WD binaries, at different stages of a continuum between intermediate and synchronised polars, suggesting a unified population of magnetic WD binaries driving coherent radio emission.
@@
White dwarf binary systems are increasingly recognised as hosts of exotic pulsar-like emission, challenging long-standing assumptions that coherent radio pulses are exclusive to neutron star systems[1, 2]. The discovery of long-period transients (LPTs), characterised by long periodicities (minutes to hours) , and strong polarisation, further deepens the mystery of how and where coherent radio emission originates [3, 4]. These intriguing systems are characterized by periodic radio pulses with high polarization (both linear and circular) and a high surface brightness pointing to a coherent radio emission process. However, the luminosities of the bright radio pulses in the $\sim$10 systems discovered thus far, exceed their rotational power, leaving uncertain the energy source. A strong magnetic field has been suggested to overcome this energy problem, suggesting at first the possibility of them being radio magnetars having perhaps had large fall-back at birth slowing down their periods [5, 6].
@@
Two LPTs, ILT J1101$+$5521 [7] and GLEAM--X J0704--37 [8, 9], have recently been associated with a WD–M dwarf binary system via the detection of the orbital motion in the optical spectrum of the M companion star, at a period coincident with the radio pulse period. These systems might be hosting a WD with a spin synchronized with the orbital motion, hence belonging to the class of polars, despite not being in an accretion phase. In contrast, the AR Scorpii (Ar Sco) and J1912–44 radio emitting binary WDs are not synchronised with their companions; coherent radio pulses are observed at beat periods resulting from the interaction between the white dwarf spin and the orbital motion. The radio emission in these systems has been intriguing all along, and scenarios involving the dipolar losses of the white dwarf rotating field and the interaction with the M–dwarf companion wind have been proposed [10, 11, 12]
@@
The phenomenological similarities between these sources invite a unified physical framework. We propose a geometric emission model in which radiation is generated when a WD’s magnetic pole sweeps across its companion star, injecting relativistic particles from the M dwarf’s wind into the WD’s magnetosphere [13, 14]. These particles are accelerated along magnetic field lines, producing a collimated radio beam aligned with the magnetic moment of the WD. The model incorporates orbital inclination, magnetic obliquity, and beam geometry, and predicts observable intensity variations as a function of the spin and orbital phase. Further intensity variations observed in the long-term radio outbursts of LPTs might be ascribed to the intrinsic variability of the M-dwarf wind.
Our model assumes a circular orbit with inclination angle $i$ between the orbital axis and the WD spin axis. The magnetic moment is tilted by angle $\alpha$ from the spin axis, and maximum radio emission is generated in the direction of the magnetic moment. Emission is modulated in two ways: (1) by the angle $\beta_{\text{MD}}$ between the magnetic pole and the M dwarf, governing the availability of particles via an assumed Gaussian function; and (2) by the angle $\beta$ between the magnetic pole and the Earth, determining visibility to observers. This model, although simplified, provides predictive power: emission is expected when both $\beta_{\text{MD}}$ and $\beta$ are small. The resulting phase-space of emission naturally leads to double-peaked pulse groups and intermittent visibility consistent with observations.
@@
GPM J1839--10 is the longest lived LPT known, exhibiting a 22-minute period with sporadic pulses observed across 36 years of archival radio data [4]. In a coordinated campaign using MeerKAT, ASKAP, and VLA in 2024, we identified a higher-order periodicity: pulses cluster in double groups occurring every $\sim$8.75 hr. We interpret this as the orbital period $T=31482.4\pm 0.2$ s, which we are able to measure to such low uncertainty by taking advantage of the 36 year baseline.
@@
Folding the data over this 8.75 hr modulation period and the previously reported 22-minute period ($P_{A}$)[4], as well as the derived beat period ($P_{B}$), reveals in Figure 1 that the pulse substructure aligns between pulses with $P_{B}=1265.2197\pm 0.0002$ s, rather than $P_{A}=1318.1957\pm 0.0002$ s. The substructure timescale of $\sim$50 s is consistent throughout the archival data, first discovered by autocorrelation between pulses [4], and folding on $P_{B}$ reveals further correlation and vertical alignment of the brightest peaks. This suggests that $P_{B}$ corresponds to the true spin period of the white dwarf (while $P_{A}$ is a spin-orbit beat period) inviting its interpretation as the cross-section of the radio beam.
The phase difference between the left and right pulse groups cannot be attributed to the difference in light travel time between orbital phases because that would be on the order of seconds for an orbit of this size, not minutes. Instead, we interpret the groups as arising from the intersections of the M dwarf’s highly inclined orbit with the cone traced out by the WD’s magnetic pole. This explains both the intermittency and the phase delay between pulse groups. Our geometric model reproduces the observed dynamic pulse profile as shown in Figure 2 when fitted using a Markov chain Monte Carlo method, with best-fit parameters indicating a near-perpendicular orbital inclination $i=100.1\pm 0.6^{\circ}$. The modelled pulse groups are the regions where the $\beta_{\text{MD}}$ contours cut into the central column of the $\beta$ colour map — where the Earth, MD, and WD pole are sufficiently aligned. No good fit exists for $P_{A}$ as the spin period.
@@
Figure 2: Modelled dynamic pulse profile of GPM J1839--10. At top left, the flux density predicted by the model $I_{\text{pred}}$ using the best-fit parameters found using MCMC is overlaid on the real pulse profiles. At top right, the colourmap is the LOS-beam angle $\beta$ as a function of spin-orbit phase, and the contours are the beam-MD angle $\beta_{\text{MD}}$. $I_{\text{pred}}$ is greatest where $\beta$ and $\beta_{\text{MD}}$ are the smallest. The bottom panel is a full orbit recorded by ASKAP and the associated model prediction. The vertical lines are spaced by the spin period.
The emission is consistently highly linearly polarised with occasional 90 jumps in polarisation angle (PA) to orthogonal polarisation modes (OPMs)[4]. However, Figure 1 reveals that the two pulse groups which appear at different orbital and spin phases are dominated by different OPMs, with a similar amount of variation in each group ($\sigma_{\text{PA}}\approx{\sim}10^{\circ}$). The PAs of individual pulses, however, may evolve more gradually and systematically across spin phase [15] than can be seen in the point clusters in the figure. It can further be seen that there is a small downturn in the PAs at around spin phase $-0.05$ (the trailing edge of the first group), as well as a similar upturn in the PAs at spin phase $0.05$ (the leading edge of the second group). Despite the possibility that pulses arriving between the two groups could potentially connect the PAs smoothly across intermediate spin/orbital phases, the flatness of the PAs of the two groups themselves, as well as the occasional $90^{\circ}$ jump within each group, lead us to prefer an OPM interpretation over a rotating vector model one[16].
Although not well understood, OPMs in pulsars are thought to be magnetospheric in origin, possibly due to birefringent properties of the co-rotating plasma[17]. We advance a similarly broad interpretation here: the OPMs are dominated by the average magnetic properties of the plasma through which the WD beam propagates along our line of sight. The consistency with which each pulse group exhibits the same dominant OPM support our geometric model, with the line of sight intersecting different physical sections of the intrabinary magnetic field during the two pulse groups. Both the single pulse PA variations and the occasional OPM jump even within a pulse group may be due to turbulence in the intrabinary wind.
Other spectral characteristics (dispersion measure, rotation measure, spectral shape) remain constant across the orbit, suggesting that variability arises primarily from geometric effects rather than intrinsic source changes on a short timescale, or propagation effects. On longer timescales (a few orbits) variability is primarily driven by MD wind activity.
To describe the system geometry, we place the WD at the origin with $\hat{z}$ as the WD spin axis, and the orbital axis inclined about $\hat{y}$. Fitting the model for GPM J1839--10, we find an orbital inclination $i=100.1\pm 0.6^{\circ}$ and magnetic obliquity $\alpha=52.1\pm 0.4^{\circ}$ with respect to $\hat{z}$. Our line of sight has azimuth $\phi_{0}=179.29\pm 0.04^{\circ}$ measured from $\hat{x}$ and zenith $\zeta=61\pm 2^{\circ}$ measured from $\hat{z}$. The radio beam has an opening angle of $W_{\text{spin}}=65\pm 3^{\circ}$ and becomes active when the MD is within $W_{\text{orb}}=70\pm 1^{\circ}$ of the WD magnetic moment $\hat{\mu}$. In Figure 3 we show the derived beam angle and orbit of the system together with critical radii to understand the emission state of the binary. The derived orbital period and WD spin period imply that: 1) the whole system is inside the WD light cylinder, 2) the companion star, if assumed to be an M-dwarf (as for the other two LPTs), is not filling its Roche-Lobe, and 3) depending on the assumed M-dwarf mass loss rate ($\sim 10^{-14}$ M yr-1 [18]), the WD Alfvén radius might be inside the corotation radius, implying that during stellar flares the system may undergo accretion episodes (possibly showing transient X-ray emission as observed in the LPT ASKAP J1832-0911 [19]), while being in propeller most of the time [20, 14]. Furthermore, the system geometry clearly shows how the rotating WD magnetosphere is naturally engulfed by the MD wind, possibly being the seed particles that are accelerated by the sweeping magnetic field lines generating the pulsed synchrotron radio emission [21, 22, 23].
Figure 3: Diagram of the binary system, in a moving reference frame centred on the WD. At left is a not-to-scale diagram of the geometric parameters. In green is WD with its magnetic moment vector. In red, the MD with its orbital path. The blue vector points towards Earth. At right is a to-scale projection of GPM J1839--10 on the y-z plane, with orbital phases marked. The cone traced by $\hat{\mu}$ crosses the y-z plane at the green full and dashed lines. The shaded red, orange, and cyan regions cover the ranges of $R_{\text{MD}}$, $R_{rl}$, and $R_{c}$ respectively, given the MD mass is $M_{\text{MD}}\in[0.14,0.5]M_{\odot}$ and the WD mass is $M_{\text{WD}}\in[0.6,1.2]M_{\odot}$. The green sector and solid red line are $W_{\text{spin}}$ and $W_{\text{orb}}$ respectively. The small and large black dashed circles are the Alfvén radii for the minimum and maximum WD masses respectively, assuming a stellar mass loss rate of $\dot{M}_{\text{MD}}$ = $10^{-14}$ M yr-1 [18].
@@
To test the generality of the model, we applied it to 7.5 hours of MeerKAT observations of the WD binary pulsar J1912--44. Folding the data over the 5.3 minute spin and 4.03 hour orbital periods [2] forms the vertically aligned dynamic pulse profile in Figure 6. We used a prior of $i=59\pm 6^{\circ}$ based on the inclination angle constraints as measured in the optical band [14]. Given the prior, we find the following constraints on the geometry: $i\in[53,65]^{\circ}$, $\alpha\in[42,70]^{\circ}$, $\phi_{0}\in[-48,-17]^{\circ}$, $\zeta\in[39,73]^{\circ}$, $W_{\text{spin}}\in[20,60]^{\circ}$, and $W_{\text{orb}}\in[67,90]^{\circ}$. These results agree broadly with the earlier proposed geometry [14], but with only one pulse group the geometrical parameters are less constrained than for GPM J1839--10.
A critical component of the model is that the direction of the radio beam is locked to the WD rotating frame in the direction of the magnetic moment. If this is not the case it becomes difficult to explain the paired pulse group morphology, and it becomes unclear why the pulse substructure would be related to the spin phase. This is difficult to reconcile with models which place the emission site at or near the companion like the electron cyclotron maser emission (ECME) model [24] proposed for ILT J1101$+$5521 or the bow shock model [12] proposed for AR Sco. One scenario which allows for an emission region locked to the WD involves relativistic electrons accelerating from the companion towards the WD along the closed magnetic field lines when the companion star’s surface is heated by magnetic interaction [25, 23]. The electrons emit synchrotron radiation as they reach magnetic mirror points on their way to the WD pole [23]. This idea has been shown to reproduce the dynamic pulse profile of Ar Sco with spin and orbit modulated components similar to our model [10].

### v2

# A binary model of long period radio transients and white dwarf pulsars

@@ Abstract
Long-period radio transients (LPTs) represent a recently uncovered class of Galactic radio sources exhibiting minute-to-hour periodicities and highly polarised pulses of second-to-minute duration. Their phenomenology does not fit exactly in any other class, although it might resemble that of radio magnetars or white dwarf (WD) radio emitting binary systems. Notably, two LPTs with confirmed multi-wavelength counterparts have been identified as WD – M dwarf binaries. Meanwhile, systems such as AR Scorpii and J1912-44 exhibit short-period pulsations in hrs-tight orbits, with polarised radio emission proposed to be generated by the interaction of the WD magnetosphere with the low-mass companion wind.
in which radio emission is triggered when the magnetic axis of a rotating WD intersects its companion’s wind in the binary orbital plane. We use a 36-year timing baseline to infer the orbital period and binary geometry from radio data alone. The model naturally predicts its intermittent emission and double-pulse structure. Crucially, we show that the beat period between the spin and the orbit matches the observed pulse substructure and polarisation signatures, providing strong support for the model. Applying it to the WD pulsar J1912-44, it successfully reproduces the emission profile and geometry as well. R1.1,R3.1]Our results suggest analogous emission-site geometries in these related classes of binary system — a possibility we extend to the broader LPT / WD pulsar population.
@@ 1 Introduction
White dwarf binary systems are increasingly recognised as hosts of exotic pulsar-like emission, challenging long-standing assumptions that coherent radio pulses are exclusive to neutron star systems[1, 2]. The discovery of long-period transients (LPTs), characterised by their long periodicities (minutes to hours) compared to the bulk of the pulsar population, further deepens the mystery of how and where pulsed coherent radio emission originates [3, 4]. These intriguing systems are characterised by periodic radio pulses with high polarisation (both linear and circular) and a high surface brightness, pointing to a coherent radio emission process. However, the luminosities of the bright radio pulses in R1.4]many of the $\sim$12 systems discovered thus far exceed their spin-down power given typical pulsar-like assumptions, leaving uncertain the energy source. A strong magnetic field has been suggested to overcome this energy problem, suggesting at first the possibility of them being radio magnetars having perhaps had large fall-back at birth slowing down their periods [5, 6].
@@
Two LPTs, ILT J1101++55 [7] and GLEAM-X J0704-37 [8, 9], have recently been associated with WD – M dwarf (MD) binary systems via the detection of the orbital motion in the optical spectrum of the M companion star, at a period coincident with the radio pulse period. These two systems might be hosting a WD with a spin synchronised with the orbital motion, despite not being in an accretion phase. R1.2,R1.5]We note, however, that the radio period need not necessarily equal the WD spin period. , which are not synchronised with their companions; coherent radio pulses are observed at the WD spin period and/or beat periods resulting from the interaction between the WD spin and the orbital motion, modulated on the orbital period with the brightest pulses arriving at a particular orbital phase. The radio emission in these systems has been intriguing all along, and scenarios involving the dipolar losses of the WD rotating field and the interaction with the MD companion wind have been proposed [12, 13, 14, 15, 16].
GPM J1839-10 is the longest-lived LPT known, exhibiting a 22-minute period with sporadic pulses observed across 36 years of archival radio data [4]. R1.8]The 30 to 300 s duration broadband radio pulses exhibit both linear and circular polarisation and quasiperiodic substructure. R1.4]Its original classification as a neutron star was based on its coherent radio luminosity ($\sim 3.5\times 10^{31}$ erg s-1 peak isotropic luminosity) but the period derivative $\dot{P}\lesssim 3.6\times 10^{-13}$ s s-1 [4] puts it in tension with the most generous pulsar emission models [18]. GPM J1839-10 also displays short duration ($\sim$20 ms) narrow-band sub-pulse features[19], suggesting multiple processes.
@@
Motivated by the detection of a longer periodicity in observations of GPM J1839-10, we put forward that it is also a WD with a low-mass stellar companion, and propose a geometric model in which radio emission from the WD polar region is modulated by the relative alignment between the WD magnetic moment and the companion star’s position. We proceed on the premise that the aforementioned alignment determines the abundance of particles injected from the companion’s wind into the WD’s magnetosphere [20, 10]. These particles are accelerated along magnetic field lines, producing a collimated radio beam aligned with the magnetic moment of the WD. The model incorporates orbital inclination, magnetic obliquity, and beam geometry, and predicts observable intensity variations as a function of the spin and orbital phase. Further intensity variations observed in the long-term radio outbursts of LPTs might be ascribed to the intrinsic variability of the companion’s wind.
R1.1,R2.1,R3.1]Similar models have been proposed for J1912-44 [10] and AR Sco [12, 13], the latter of which reproduces AR Sco’s more complex pulse morphology with spin and orbit modulated components resembling GPM J1839-10. The phenomenological similarities between these sources and the geometric similarity invite an analogous physical framework.
2 Results
@@
In a coordinated campaign using MeerKAT, ASKAP, and VLA in 2024, we obtained a near-contiguous 36 hour track of GPM J1839-10, in which we identified a higher-order periodicity: pulses cluster in paired groups occurring every $\sim$8.75 hr. We interpret this as an orbital period $P_{\text{orb}}=31482.4\pm 0.2$ s, which we are able to measure to such low uncertainty by taking advantage of the 36 year baseline (see subsection 4.3).
@@
Folding the data over this 8.75 hr modulation period and the previously reported[4] 22-minute period ($P_{1}$), as well as the beat period ($P_{2}$) between $P_{1}$ and $P_{\text{orb}}$, creates the dynamic pulse profiles in Figure 1, revealing that the pulse substructure aligns between pulses with $P_{2}=1265.2197\pm 0.0002$ s, rather than $P_{1}=1318.1957\pm 0.0002$ s. The substructure timescale of $\sim$50 s is consistent throughout the archival data, first discovered by autocorrelation between pulses [4], and folding on $P_{2}$ reveals further correlation and vertical alignment of the brightest peaks. This suggests that $P_{2}$ corresponds to the true spin period of the white dwarf (while $P_{1}$ is a spin-orbit beat period) inviting its interpretation as the cross-section of the radio beam.
@@
2.1 Modelling
The phase difference between the left and right pulse groups in Figure 1 cannot be attributed to the difference in light travel time between orbital phases because that would be on the order of seconds for an orbit of this size, not minutes. Instead, we consider a WD with a magnetic moment $\vec{\mu}$ misaligned from its spin axis $\hat{z}$ by angle $\alpha$ and a companion low-mass star (assumed to be an MD for simplicity) in a circular orbit whose axis is inclined by angle $i$ from $\hat{z}$ about $\hat{y}$ (see Figure 3). When the spin and orbital axes are misaligned, there are two orbital phases at which the rotating magnetic moment vector intersects the MD, which we suggest correspond to the observed pulse groups. Therefore, we model the observed radio emission as modulated by
  1. 1.
    a Gaussian function of the angle $\beta_{\text{MD}}$ between the WD magnetic moment and the MD, governing the availability of the MD’s wind at the WD pole, and
  2. 2.
    a Gaussian function of the angle $\beta$ between the WD magnetic moment and the line of sight (LOS), which is the cross-section of the radio beam parallel to the WD magnetic moment.
This model, although simplified, provides predictive power: emission is expected when both $\beta_{\text{MD}}$ and $\beta$ are small. The resulting phase-space of emission naturally leads to double-peaked pulse groups and intermittent visibility consistent with observations.
Fitting the model for GPM J1839-10 using a Markov chain Monte Carlo method, we find a near-perpendicular orbital inclination $i=100.1\pm 0.6^{\circ}$ and magnetic obliquity $\alpha=52.1\pm 0.4^{\circ}$. Our LOS has azimuth $\phi_{0}=179.29\pm 0.04^{\circ}$ and zenith $\zeta=61\pm 2^{\circ}$. The radio beam has an opening angle of $W_{\text{spin}}=65\pm 3^{\circ}$ and becomes active when the MD is within $W_{\text{orb}}=70\pm 1^{\circ}$ of $\hat{\mu}$. Figure 2 is the predicted dynamic pulse profile, alongside a map of $\beta$ and $\beta_{\text{MD}}$, and the lightcurve of a representative orbit. Intuitively, the modelled pulse groups are the regions where the $\beta_{\text{MD}}$ contours cut into the central column of the $\beta$ colour map — where the Earth, MD, and WD pole are sufficiently aligned. The model also predicts weak emission in the central spin phase $\sim$ orbit phase $\sim$ 0 region, where we indeed see sporadic dim pulses at about the substructure timescale. No good fit exists for $P_{1}$ as the spin period. In Figure 3 we show the derived beam angle and orbit of the system together with critical radii on a 2D projection.
@@ 2.1 Modelling
To test the generality of the model, we applied it to previously published [2] 7.5 hours of MeerKAT observations of the WD binary pulsar J1912-44. Folding the data over the 5.3 minute spin and 4.03 hour orbital periods forms the vertically aligned dynamic pulse profile in Figure 4. We used a prior of $i=59\pm 6^{\circ}$ based on the inclination angle constraints as measured in the optical band [10]. Given the prior, we find the following constraints on the geometry: $i\in[53,65]^{\circ}$, $\alpha\in[42,70]^{\circ}$, $\phi_{0}\in[-48,-17]^{\circ}$, $\zeta\in[39,73]^{\circ}$, $W_{\text{spin}}\in[20,60]^{\circ}$, and $W_{\text{orb}}\in[67,90]^{\circ}$. These results agree broadly with the earlier proposed geometry [10], but with only one pulse group the geometrical parameters are significantly less constrained than for GPM J1839-10. The degeneracy of the J1912-44 parameters suggests that the vertical single-pulse-group morphology should be the most commonly observed among WD pulsars and LPTs.
Figure 2: Modelled dynamic pulse profile of GPM J1839-10. At top left, the flux density predicted by the model $I_{\text{pred}}$ using the best-fit parameters found using MCMC is overlaid on the real pulse profiles. At top right, the colourmap is the LOS-beam angle $\beta$ as a function of spin-orbit phase, and the contours are the beam-MD angle $\beta_{\text{MD}}$. The bottom panel is a full orbit recorded by ASKAP normalised to 1 GHz and the associated model prediction. The vertical lines are spaced by the spin period.
Figure 3: Diagram of the binary system, in a moving reference frame centred on the WD. At left is a not-to-scale diagram of the geometric parameters. In green is WD with its magnetic moment vector. In red, the MD with its orbital path. The blue vector points towards Earth. At right is a to-scale projection of GPM J1839-10 on the y-z plane, with orbital phases marked. The cone traced by $\hat{\mu}$ crosses the y-z plane at the green full and dashed lines. The shaded red, orange, and cyan regions cover the ranges of $R_{\text{MD}}$, $R_{rl}$, and $R_{c}$ respectively, given the MD mass is $M_{\text{MD}}\in[0.14,0.5]M_{\odot}$ and the WD mass is $M_{\text{WD}}\in[0.6,1.2]M_{\odot}$. The green sector and solid red line are $W_{\text{spin}}$ and $W_{\text{orb}}$ respectively. The small and large black dashed circles are the Alfvén radii for the minimum and maximum WD masses respectively, assuming a stellar mass loss rate of $\dot{M}_{\text{MD}}$ = $10^{-14}$ M yr-1 [21].
Figure 4: Modelled dynamic pulse profile for J1912-44.
3 Discussion
R2.2]Strengthening the case for a WD LPT progenitor beyond the thus-far mentioned are the LPTs ASKAP J1448-68 [22] and CHIME/ILT J1634++44 [23, 24], both of which have been shown to have optical spectral properties consistent with a WD. Furthermore, hints of accretion in the former and pulse arrival time properties in the latter imply possible binary systems.
For GPM J1839-10, Figure 3 shows how the rotating WD magnetosphere is naturally engulfed by the MD wind, possibly being the seed particles that are accelerated by the magnetic field lines generating the pulsed radio emission [25]. It further shows that: 1) the whole system is inside the WD light cylinder, 2) the companion star, if assumed to be an MD (as for the other two LPTs), is not filling its Roche-Lobe, and 3) depending on the assumed M-dwarf mass loss rate ($\sim 10^{-14}$ M yr-1 [21]), the WD Alfvén radius might be inside the corotation radius, implying that during stellar flares the system may undergo accretion episodes (possibly showing transient X-ray emission as observed in the LPT ASKAP J1832-0911 [26], R1.4,R3.6]although the X-ray luminosity limit of $L_{X}<2\times 10^{33}$ erg s-1[4] for GPM J1839-10 does not preclude some level of accretion[27]), while being in propeller most of the time [28, 10]. ASKAP J1448-68 has also been observed to emit X-rays[22], and its multi-wavelength properties are consistent with a magnetic white dwarf, possibly an accreting binary system. In general, it seems that significant accretion may not be a necessary component of the radio emission mechanism (whatever it may be) since we both do [26, 22, 23, 24, 2] and do not [7, 8, 1, 17] see signs of such across the LPT / WD pulsar population.
1.8,R2.1,3.1]It has been shown [12, 13] that for AR Sco’s radio emission, like for that of GPM J1839-10 in this work, the pulse morphology is best modelled with a beam direction locked to the WD rotating frame, as opposed to the binary frame, and modulated by the companion’s location in that frame. One may expect the geometric features between the discussed LPTs and WD pulsars to be similar simply because both originate from binary systems, but the comparison is stronger than that. These geometric constraints give us the important clue that the emission direction is dominated by the WD’s local field, while powered by interaction with the MD. If this is not the case it becomes difficult to explain the paired pulse group morphology, and it becomes unclear why the pulse substructure would be related to the spin phase. R1.9]This is difficult to reconcile with models like electron cyclotron maser emission (ECME) [29] or models which place the emission site near the MD [15, 30].
A scenario which has been suggested for AR Sco [1, 25] which allows for an emission region locked to the WD involves relativistic electrons accelerating from the companion towards the WD along the closed magnetic field lines when the companion star’s surface is heated by magnetic interaction. The electrons emit synchrotron radiation as they reach magnetic mirror points on their way to the WD pole. This idea was used successfully in the earlier mentioned geometric model [12] to predict polarimetric properties of AR Sco. R1.8]The broadband nature and spectral shape of GPM J1839-10 could be consistent with this, but the radio luminosity is orders of magnitude brighter. Despite the possibility that the radio emission from AR Sco and GPM J1839-10 are powered by different mechanisms, we highlight the consistent interpretation in the two systems that the emission is modulated as the source of particles (i.e. the MD) moves along its orbit. If the mechanisms differ then it is likely in the acceleration mechanism of said particles.
R1.9]Alternative forms of ECME in LPTs have been proposed and remain viable, such as relativistic ECME [14] or loss-cone-driven maser (LCDM) [16], which produce beamed emission aligned with the local magnetic field of the WD. R3.5]Less luminous ($L\lesssim 10^{25}$ erg s-1) radio emission from a related class of WD binaries, (magnetic) cataclysmic variables (CV / mCV), has been observed[31, 32]. If this emission is ECME, consistent with their typical circular polarisation [33] (not a requirement in the relativistic case), then there could be a natural spectrum from non-relativistic to relativistic regimes.
R1.6,R1.9,R3.2]In the case of ILT J1101++55, relativistic ECME was proposed to be powered by unipolar induction, in which an electric field is induced in the MD by the penetrating WD magnetic field, leading to the establishment of a current loop along the field lines connecting the two. However, this requires that the orbital separation is sufficiently small that the WD magnetic field dominates at the location of the MD. According to recent theoretical work [16], when the orbit is greater than $P_{\text{orb}}\sim 3$ hr, the MD magnetic field is strong enough to prevent the WD field from penetrating its surface, and we transition from the unipolar induction phase to the magnetospheric interaction phase. In this phase, radio emission occurs along magnetic field lines connecting the WD to a magnetic reconnection region between the two, rather than the MD itself. The orbit of GPM J1839-10 would place it in the magnetospheric interaction phase, and we calculate that the observed luminosity can be reached without requiring a strongly magnetic WD.